OpenStudy (anonymous):

In triangle ABC, a = 3, b = 5, and c = 7. Find the approximate value of angle A.

OpenStudy (mathmate):

Hint: cosine rule!

OpenStudy (anonymous):

oh ok use the quadratic formula

OpenStudy (anonymous):

22° 38° 142° 158° these are the answer choices

OpenStudy (anonymous):

\[\frac{ -b+-\sqrt{(b)-4(a)(c)} }{ ?2(a) }\]

OpenStudy (anonymous):

plug it in

OpenStudy (anonymous):

ok one second

OpenStudy (anonymous):

im still not understanding

OpenStudy (mathmate):

In geometry, it always helps to draw a diagram according to the given information. 

OpenStudy (mathmate):

cosine rule says: \(a^2=b^2+c^2-2(b)(c) cos(A)\) from which you can solve for cos(A): \(\Large cos(A)=\frac{b^2+c^2-a^2}{2bc}\) So you can substitute a,b,c into the equation and solve for angle A.

OpenStudy (anonymous):

and what do you get when you that because i keepp getting something different

OpenStudy (anonymous):

@kyrabaaker

OpenStudy (mathmate):

@jcwilliams504 What have you done so far?

OpenStudy (anonymous):

i plugged in everything but i dont know how to solve

OpenStudy (mathmate):

Do you know the values of a, b, and c?

OpenStudy (anonymous):

3,5, and 7

OpenStudy (mathmate):

Good, so what did you get for: \(\Large \frac{b^2+c^2-a^2}{2bc}\)

OpenStudy (anonymous):

65/70 :/

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

That's correct. You can find the angle A by solving cos(A)=65/70=13/14 or A = cos\(^{-1}\)(13/14)

OpenStudy (anonymous):

whats after that?